# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

### v1

@@ Abstract
Robust verbal confidence generated by large language models (LLMs) is crucial for the deployment of LLMs to ensure transparency, trust, and safety in human-AI interactions across many high-stakes applications. In this paper, we present the first comprehensive study on the robustness of verbal confidence under adversarial attacks. We introduce a novel framework for attacking verbal confidence scores through both perturbation and jailbreak-based methods, and show that these attacks can significantly jeopardize verbal confidence estimates and lead to frequent answer changes. We examine a variety of prompting strategies, model sizes, and application domains, revealing that current confidence elicitation methods are vulnerable and that commonly used defence techniques are largely ineffective or counterproductive. Our findings underscore the urgent need to design more robust mechanisms for confidence expression in LLMs, as even subtle semantic-preserving modifications can lead to misleading confidence in responses.
@@ 1 Introduction
With the unprecedented advancement of large language models (LLMs) [1, 42, 45] and their widespread deployment in real-world systems, the robustness in acquiring LLMs’ confidence is crucial for a wide range of applications [7, 14, 66, 71] including high-stakes tasks [44, 67]. For human beings, the expression of epistemic uncertainty using natural language is an innate ability. We envision in the near future it will be crucial for AI systems to express their confidence verbally and do so accurately to produce seamless interactions in human-AI collaborations.
@@
The ability of LLMs to verbally convey their confidence [15], called verbal confidence in this paper, is not only important for applications, but also an indicator of the intelligence of LLMs. Ensuring accurate and robust verbal confidence also underlies the objective of aligning AI models’ core values with those of humans, such as honesty [34, 69], which could be crucial to avoiding mistrust [9, 53, 72] and maintaining future AI to be safe.
@@
Figure 1: (a) Demonstration of how perturbations on prompts can affect final verbal confidence (b) An overview of our proposed attack framework centered on using generated confidence scores to optimize attacks that either perturb the prompt or append a series of optimized trigger tokens that ultimately result in a reduction in confidence.
In addition, many state-of-the-art (SOTA) LLMs do not allow direct access to model information which remains a crucial limitation for uncertainty estimation [67, 71]. Verbal confidence enables the estimation of uncertainty strictly using generated outputs, leading to a considerable amout of recent research on examining the role of verbal confidence [14, 20, 30, 35, 41, 51, 58, 59, 66, 73] as well as the influence and implication of verbal confidence on trustworthiness and transparency [24]. Real-world industry systems have also utilized verbal confidence [39, 71, 7], where analyzing LLMs’ behaviour on this front is key for successful operation.
@@
Our work presents the first comprehensive study on the robustness of verbal confidence under adversarial attacks and addresses the basic question: how robust is the verbal confidence of the existing large language models? We highlight the differences in how models perceive and manage uncertainty cues compared to how humans naturally understand them. Furthermore, given the ease of attaining verbal confidence from even the most guarded black-box models, it is key to determine how effectively these can be leveraged by adversaries as an attack objective to alter model behaviour in other ways, such as generated answers.
@@
To assess the Achilles’ heels of LLMs models in expressing their confidence [5, 52, 54, 65, 75], we propose attack models designed for attacking verbal confidence scores through both perturbation and jailbreak-based approaches, and show that these attacks can significantly jeopardize verbal confidence estimates. We explore three key questions for model deployment that have not been explored: (i) How vulnerable are these estimates under a range of adversarial attack techniques? (ii) How can these attacks be effectively crafted to target confidence and answer changes? (iii) How easily can such attacks be defended against with existing defending methods? Detailed discussions on real-world applications and the importance of maintaining proper confidence can be found in Appendix A.
In brief, the main contributions of our work are summarized as follows:
@@
  • We design novel verbal confidence attacks based on both perturbation-based and jailbreak-based approaches, showing they can cause up to a 30% reduction in average confidence on user queries and prompt a high rate of answer changes (up to 100% in originally correctly predicted samples).
@@ 2 Related Work
Adversarial Attacks in NLP. Adversarial attacks have served as a critical tool to reveal key vulnerabilities of vision and NLP models [18, 25, 52]. Generally adversarial attack methods can be divided into those that are perturbation-based and those that are jailbreak-based. Recent works of the former type have examined the robustness of LLMs under word level attacks that target prediction and generation accuracy  [62, 74] through threat vectors such as demonstrations [63]. Jailbreak-based attack methods involve manually or algorithmically [5, 65] generated trigger phrases that are appended to the input prompt to bypass safety guardrails [8]. The first methods utilized means such as optimized universal adversarial triggers (UAT) [60] and top-$k$ gradient-based search [56]. Recent variants of trigger attacks include the GCG attack [75], black-box only approaches [32], and those with different optimization methods such as genetic algorithms [37], policy gradient methods [10], and using specially trained LLMs[48].
@@
Compared to attacks that targeting victim models’ accuracy, adversarial attacks against models’ confidence has been understudied but has become more prominent as AI models are deployed into more applications where calibration robustness needs to investigated (see further discussion in Appendix A). The previous works targeted logits-based confidence, where confidence is obtained from neural networks themselves. For example, Galil and El-Yaniv [13] formulated confidence attacks and Emde et al. [11] examined confidence attacks through certified robustness guarantees. Obadinma et al. [43] designed a framework that tunes adversarial attacks to target miscalibration. Zeng et al. [70] examined attacks that target uncertainties derived from output tokens’ logits. None of these approaches, however, have investigated verbal confidences, which, as discussed above, is a basic problem that needs to be further understood.
@@
Recently, a significant amount of research has been conducted to examine the effectiveness of explicitly generating confidence scores alongside answers (see Geng et al. [15] for a detailed survey). For example, Lin et al. [35] examined fine-tuning decoder-only LLMs to generate calibrated confidence. Tian et al. [59] evaluated prompting models for confidence scores and gauge the calibration accordingly. Xiong et al. [66] examined multiple prompting, sampling, and aggregation strategies for verbal confidence scores. Our research is the first to examine the robustness of verbal confidence through adversarial attacks.
@@ 3 Adversarial Attacks on Verbal Confidence
We investigate the robustness of verbal confidence produced by LLMs. Given a user prompt consisting of a query $\bm{\mathbf{X}}\equiv\{x_{1},\dots,x_{N}\}$ with a length of $N$ word tokens, users of an LLM can obtain a confidence level by providing the LLM with an additional task prompt $\bm{\mathcal{P}}\equiv\{p_{1},\dots,p_{M}\}$ of $M$ word tokens, which describes the manner in which the LLM should provide a confidence score and gives an illustrative demonstration of how to do it (i.e., through an example). The queried LLM then generates a sequence of tokens $\bm{\mathcal{Y}}\equiv\{y_{1},\ldots,y_{R}\}$ consisting of $R$ word tokens as a response, containing a subsequence representing its overall answer $\mathcal{\bm{A}}\subseteq\bm{\mathcal{Y}}$ to the query. A corresponding confidence score, which we call the verbal confidence $\mathcal{C}$, is a scalar quantity attained with the help of the task-specific prompt that reflects the model’s belief in the correctness of its response $\mathcal{A}$. The score $\mathcal{C}$ is derived strictly from the output word tokens $\bm{\mathcal{Y}}$ alone without any knowledge of internal model states. We refer to the entire process as confidence elicitation (CE). Formally, confidence elicitation methods (CEMs) are functions that map a generated sequence of word tokens $\bm{\mathcal{Y}}$ to a confidence score, $\mbox{CEM}:\bm{\mathcal{Y}}\rightarrow\mathcal{C}$, given a specific LLM that serves as a mapping function from the input prompts to the output sequence $\mbox{LLM}:(\mathbf{X},\bm{\mathcal{P}})\rightarrow\bm{\mathcal{Y}}$. As detailed in Section 3.4, the specific CEM functions can be implemented with different prompting strategies, including chain-of-thought, multi-step, and self-consistency prompting.
@@ 3.2 Perturbation-based Confidence Attack
We propose two perturbation-based verbal confidence attacks (VCA): VCA-TF and VCA-TB. These VCA attacks are based on two popular black-box attack algorithms: Textfooler (TF) [28] and Textbugger (TB) [33]. However, unlike the original algorithms, we design new scoring functions which are optimized using verbal confidence in contrast to class-wise probabilities derived from model logits. The key principal behind these methods is to generate importance scores for each token in the sequence by determining the corresponding difference in the predicted confidence without the token present in the sequence, before ranking each token and prioritizing perturbations on tokens with the highest difference. VCA-TF focuses on synonym substitutions, while VCA-TB includes character modifications like common typos.
@@
5:    $\mathcal{L}_{z}$leftarrow AverageConfidence($z$, $S$) $\triangleright$ w/o any trigger
@@
12:           $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) -- $\mathcal{L}_{z}$
@@
19:    while $|\mathbf{P}_{g}|$ $<$ $|\mathbf{P}|$  do
20:        $p_{ADV_{1}}$, $p_{ADV_{2}}$leftarrowTournamentSelect($K=2$, $\mathbf{P}$)
@@
23:        $p_{ADV_{1}}$, $p_{ADV_{2}}$leftarrow Cross-Over($p_{ADV_{1}}$,$p_{ADV_{2}}$, $v$)
@@
29:    $\mathbf{P}$leftarrow $\mathbf{P}_{g}$
@@
38:$\mathbf{T}$leftarrow $\mathbf{P}[argmin(\mathcal{L}_{p_{G}})]$triangleright Prompt with lowest loss
@@ 3.3.1 ConfidenceTriggers
In Figure 1(b), we show the potential for a more generalizable form of attack by designing a VCA method called ConfidenceTriggers, which is designed to target confidence based on a popular jailbreak approach [32]. By optimizing a single series of trigger tokens that get appended to any number of prompts, LLMs are led to generate lowered confidences scores for these prompts. Note that our work is the first to optimize and study triggers that target verbal confidence scores.
@@
Triggers are optimized using a completely black-box genetic algorithm (Algorithm 1). No model-specific information or tokenizers are required, corresponding to the setup that needs verbal confidence. To optimize for confidence reduction, we first generate an initial set of $\alpha$ prompts consisting of $\beta$ number of random words. These words come from a list of approximately 2000 words related to the topic of uncertainty to create an initial focused search space. This initialization process is accomplished using the InitializePrompts function in the algorithm. We append the triggers to the system prompt, as we will later show it is a more effective attack vector. This also has the byproduct of if a system or API is infiltrated with such a trigger, then confidence estimates on subsequent (benign) user queries are compromised. The loss function to be minimized is the difference in average verbal confidence (see AverageConfidence function) across $\xi$ randomly sampled training examples when using the tested trigger compared to the average confidence estimated on $S$ randomly sampled examples without the trigger.
@@
To enable generalizability across different CEMs, we use four different variants to estimate the loss for each trigger. To produce prompts for the next generation, first $E$ elites are selected, then pairs of parent triggers (from the previous generation) are iteratively selected through top-$k$ tournament selection [4] with $k=2$ to undergo Single Point Cross-Over [23] and Random (Uniform) Mutation [17] until there are $N$ child prompts. For Cross-Over, a random split point is chosen and two child prompts are created whereby the first consists of the sequence from the first parent up to the split point, and then from the second parent after the split point, and vice versa for the second child. For Mutation, a third child is created by iteratively sampling token randomly from either the first parent or second parent. 20% of indices are chosen as mutation points, where a random word is sampled instead. After $G$ generations, the lowest loss prompt is selected. The efficiency of this attack algorithm is dependent on the input parameters in Algorithm 1, with the equation for number of model calls needed to optimize a set of triggers being approximately $|\mathbf{Z}|*S+G*\xi*|Z|*\alpha+\alpha*|\mathbf{Z}|*S$, although once optimized a set of tokens can be reused indefinitely. Detailed descriptions of settings and the Cross-Over and Mutation functions can be found in Appendix C.1.

### v3

@@ Abstract
Robust verbal confidence generated by large language models (LLMs) is crucial for the deployment of LLMs to help ensure transparency, trust, and safety in many applications, including those involving human-AI interactions. In this paper, we present the first comprehensive study on the robustness of verbal confidence under adversarial attacks. We introduce attack frameworks targeting verbal confidence scores through both perturbation and jailbreak-based methods, and demonstrate that these attacks can significantly impair verbal confidence estimates and lead to frequent answer changes. We examine a variety of prompting strategies, model sizes, and application domains, revealing that current verbal confidence is vulnerable and that commonly used defence techniques are largely ineffective or counterproductive. Our findings underscore the need to design robust mechanisms for confidence expression in LLMs, as even subtle semantic-preserving modifications can lead to misleading confidence in responses.
@@ 1 Introduction
With the unprecedented advancement of large language models (LLMs) [llama3modelcard, minaee2024largelanguagemodelssurvey, OpenAI2024] and their widespread deployment in real-world systems, the robustness in acquiring LLMs’ confidence is crucial for a wide range of applications [chen-mueller-2024-quantifying, Geng2023ASO, xiong2024can, zhang2023pacelmpromptingaugmentationcalibrated] including high-stakes tasks [Omar2024.08.11.24311810, XU2025103455]. For human beings, the expression of epistemic uncertainty using natural language is an inherent ability. As the intelligence of AI systems advances rapidly, we envision in the future it will be important for AI systems to express their confidence verbally and accurately to produce seamless interactions in human-AI collaborations.
@@
The ability of LLMs to verbally convey their confidence [geng2024survey], referred to as verbal confidence in this paper, is not only important for applications, but also an indicator of the intelligence of LLMs. Ensuring accurate and robust verbal confidence also underlies the objective of aligning AI models’ core values with those of humans, such as honesty [li2024survey, Yang2023AlignmentFH], which could be critical to avoiding mistrust [dhuliawala-etal-2023-diachronic, 10.1145/3491102.3501967, 10.1145/3351095.3372852] and maintaining future AI to be safe.
Figure 1: (a) Demonstration of how perturbations on prompts can affect final verbal confidence
@@
(b) An overview of our proposed attack framework centered on using generated confidence scores to optimize attacks that either perturb the prompt or append a series of optimized trigger tokens that ultimately result in a reduction in confidence.
Many state-of-the-art (SOTA) LLMs do not allow direct access to model information which remains an essential limitation for confidence estimation [XU2025103455, zhang2023pacelmpromptingaugmentationcalibrated]. Even if they are available, logit-based confidences centered on attaining tokens probabilities and the aggregation of such scores may not reflect the true model confidence on the overall answer. Verbal confidence encourages AI systems to express their estimation of uncertainty and has attracted a considerable amount of recent research, which examine the mechanism and impact of LLMs’ verbal confidence [Geng2023ASO, groot-valdenegro-toro-2024-overconfidence, kumar-etal-2024-confidence, lin2022teaching, mielke-etal-2022-reducing, portillowightman2023strength, sun2024confidenceestimationllmbaseddialogue, DBLP:journals/corr/abs-2305-14975, xiong2024can, zhou-etal-2023-navigating], including the influence and implications on trustworthiness and transparency [huang2024trustllmtrustworthinesslargelanguage]. Numerous real-world industry systems have started to utilize verbal confidence [chen-mueller-2024-quantifying, macdonald2024generative, zhang2023pacelmpromptingaugmentationcalibrated], motivating the analysis of LLMs’ behaviour on this front as it is key for their successful operation.
@@
Our work presents the first comprehensive study on the robustness of verbal confidence under adversarial attacks and addresses the basic question: how robust is the verbal confidence of existing large language models? We highlight the differences in how models perceive and manage uncertainty cues compared to how humans naturally understand them. Furthermore, given the ease of attaining verbal confidence from even the most guarded black-box models, it is key to determine how effectively these can be leveraged by adversaries as an attack objective to alter model behaviour in other ways, such as predicted answers.
@@
To assess the Achilles’ heel of LLMs in expressing their confidence [carlini2023are, QIU2022278, REN2020346, wei2023jailbroken, zou2023universal], we propose a framework designed for attacking verbal confidence scores through both perturbation and jailbreak-based approaches, and show that these attacks can significantly jeopardize verbal confidence estimates. We explore three key questions for model deployment that have not been explored: (i) How vulnerable are these estimates under a range of adversarial attack techniques? (ii) How can these attacks be effectively crafted to target confidence and answer changes? (iii) How easily can such attacks be detected or neutralized with existing defence methods? Detailed discussions on real-world applications and the importance of maintaining proper confidence can be found in Appendix A.
In summary, the main contributions of our work are as follows:
@@
  • We design novel verbal confidence attacks based on both perturbation-based and jailbreak-based approaches, showing they can cause up to a 30% reduction in average confidence on user queries and induce a high rate of answer changes (up to 100% in originally correctly predicted samples).
@@ 2 Related Work
Adversarial Attacks in NLP. Adversarial attacks have served as a critical tool to reveal key vulnerabilities of vision and NLP models [10.1145/3593042, huq2020adversarial, QIU2022278]. Adversarial attack methods can be divided into perturbation-based and jailbreak-based attacks. Recent works of the former type have examined the robustness of LLMs under word-level attacks that target prediction and generation accuracy  [wang2023on, zhu2023promptbench] through threat vectors such as demonstrations [wang2023adversarial]. Jailbreak-based attacks involve manually or algorithmically [carlini2023are, wei2023jailbroken] generated trigger phrases that are appended to the input prompt to bypass safety guardrails [Deng2023MASTERKEYAJ]. Early methods utilized means such as optimized universal adversarial triggers (UAT) [wallace-etal-2019-universal] and top-$k$ gradient-based search [shin-etal-2020-autoprompt]. Recent variants of trigger attacks include the GCG attack [zou2023universal], black-box only approaches [lapid2023open], and those with different optimization methods such as genetic algorithms [liu2024autodan], policy gradient methods [diao2023blackbox], and using specially trained LLMs [paulus2024advprompterfastadaptiveadversarial].
@@
Compared to attacks that target victim models’ accuracy, adversarial attacks against models’ confidence has been understudied but have become more prominent as AI models are deployed into more applications where calibration robustness needs to investigated (see further discussion in Appendix A). Previous works primarily targeted logit-based confidence. For example, galil2021disrupting formulated confidence attacks and emde2023certified examined confidence attacks through certified robustness guarantees. obadinma2024calibration designed a framework that tunes adversarial attacks to target miscalibration. zeng2024uncertaintyfragilemanipulatinguncertainty examined attacks that target uncertainties derived from output tokens’ logits. None of these approaches, however, have investigated the vulnerability of verbal confidence. This, as discussed above, is a fundamental problem warranting further investigation.
@@
Recently, a significant amount of research has been conducted to examine the effectiveness of explicitly generating confidence scores alongside natural language responses (see geng2024survey for a detailed survey). For example, lin2022teaching examined fine-tuning decoder-based LLMs to generate calibrated confidence. DBLP:journals/corr/abs-2305-14975 evaluated prompting models for confidence scores and gauge the calibration accordingly. xiong2024can examined multiple prompting, sampling, and aggregation strategies for verbal confidence scores. To the best of our knowledge, this research is the first to examine the robustness of verbal confidence through adversarial attacks.
@@ 3 Adversarial Attacks on Verbal Confidence
We investigate the robustness of verbal confidence produced by LLMs. Given a user prompt consisting of a query $\bm{\mathbf{X}}\equiv\{x_{1},\dots,x_{N}\}$ with a length of $N$ word tokens, users of an LLM can obtain a confidence level by providing the LLM with an additional task prompt $\bm{\mathcal{P}}\equiv\{p_{1},\dots,p_{M}\}$ of $M$ word tokens, which describes how the LLM should provide a confidence score and may give an illustrative demonstration of how to do it (i.e., through an example). The queried LLM then generates a sequence of tokens $\bm{\mathcal{Y}}\equiv\{y_{1},\ldots,y_{R}\}$ consisting of $R$ word tokens as a response, containing a subsequence representing its overall answer $\mathcal{\bm{A}}\subseteq\bm{\mathcal{Y}}$ to the query. A corresponding confidence score, which we call the verbal confidence $\mathcal{C}$, is a scalar quantity attained with the help of the task-specific prompt that reflects the model’s belief in the correctness of its answer $\mathcal{A}$. The score $\mathcal{C}$ is derived strictly from the output word tokens $\bm{\mathcal{Y}}$ alone without any knowledge of internal model states. We refer to the entire process as confidence elicitation (CE). Formally, confidence elicitation methods (CEMs) are functions that map a generated sequence of word tokens $\bm{\mathcal{Y}}$ to a confidence score, $\mbox{CEM}:\bm{\mathcal{Y}}\rightarrow\mathcal{C}$, given a specific LLM that serves as a mapping function from the input prompts to the output sequence $\mbox{LLM}:(\mathbf{X},\bm{\mathcal{P}})\rightarrow\bm{\mathcal{Y}}$. As detailed in Section 3.4, the specific CEM functions can be implemented with different prompting strategies, including Chain-of-Thought and Multi-Step prompting.
@@ 3.2 Perturbation-based Confidence Attack
We propose two perturbation-based verbal confidence attacks: VCA-TF and VCA-TB. These VCA attacks are based on two popular black-box attack algorithms: Textfooler (TF) [DBLP:conf/aaai/JinJZS20] and Textbugger (TB) [DBLP:conf/ndss/LiJDLW19]. However, unlike the original algorithms, we design new scoring functions which are optimized using verbal confidence in contrast to class-wise probabilities derived from model logits. These methods first generate importance scores for each token in the sequence by determining the corresponding difference in the predicted confidence without the token present in the sequence. Then, each token is ranked and perturbations are prioritized on tokens with the highest difference. VCA-TF focuses on synonym substitutions, while VCA-TB also includes character modifications like common typos.
@@
5:  $\mathcal{L}_{z}$ $\leftarrow$ AverageConfidence($z$, $S$) $\triangleright$ w/o any trigger
@@
12:     $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) - $\mathcal{L}_{z}$
@@
19:  while $|\mathbf{P}_{g}|$ << $|\mathbf{P}|$ do
[…]
@@
23:   $p_{ADV_{1}}$, $p_{ADV_{2}}$ $\leftarrow$ Cross-Over($p_{ADV_{1}}$$p_{ADV_{2}}$, $v$)
@@
29:  $\mathbf{P}$ $\leftarrow$ $\mathbf{P}_{g}$
@@
[…]
@@ 3.3.1 ConfidenceTriggers
In Figure 1(b), we show the potential for a more generalizable form of attack by designing a VCA method called ConfidenceTriggers, which is designed to target confidence based on a popular jailbreak approach [lapid2023open]. By optimizing a single series of trigger tokens that get appended to any number of prompts, LLMs are led to generate lowered confidences scores for these prompts. Our work is the first to optimize and study triggers that target verbal confidence scores.
@@
Triggers are optimized using a completely black-box genetic algorithm (Algorithm 1). No model-specific information or tokenizers are required, corresponding to the setup that only needs verbal confidence. To optimize for confidence reduction, we first generate an initial set of $\alpha$ prompts consisting of $\beta$ number of random words. These words come from a list of approximately 2,000 words related to the topic of uncertainty to create an initial focused search space. This initialization process is accomplished using the InitializePrompts function in the algorithm. We append the triggers to the system prompt, as we will later show it is a more effective attack vector. This also has the byproduct that if a system or API is infiltrated with such a trigger, then confidence estimates on subsequent (benign) user queries are compromised. The loss function to be minimized is the difference in average verbal confidence (see AverageConfidence function in Algorithm 5) across $\xi$ randomly sampled training examples when using the tested trigger compared to the average confidence estimated on $S$ randomly sampled examples without the trigger.
@@
To enable generalizability across different CEMs, we use four different variants to estimate the loss for each trigger (detailed below in Section 3.4). To produce prompts for the next generation, first $E$ number of elites are selected, then pairs of parent triggers (from the previous generation) are iteratively selected through top-$k$ tournament selection [10.1162/evco.1996.4.4.361], with $k=2$ to undergo Single Point Cross-Over [10.7551/mitpress/1090.001.0001] and Random (Uniform) Mutation [10.5555/534133] until there are $\alpha$ child prompts. For Cross-Over, a random split point is chosen and two child prompts are created whereby the first consists of the sequence from the first parent up to the split point, and then from the second parent after the split point, and vice versa for the second child. For Mutation, a third child is created by iteratively sampling token randomly from either the first parent or second parent. To be specific, 20% of indices are chosen as mutation points, where a random word is sampled instead. After $G$ generations, the lowest loss prompt is selected. The efficiency of this attack algorithm is dependent on the input parameters in Algorithm 1, with the equation for number of model calls needed to optimize a set of triggers being approximately $|\mathbf{Z}|*S+G*\xi*|\mathbf{Z}|*\alpha+\alpha*|\mathbf{Z}|*S$, although once optimized a set of tokens can be reused indefinitely. Detailed descriptions of settings and the Cross-Over and Mutation functions can be found in Appendix C.1.